# RDDM: Practicing RAW Domain Diffusion Model for Real-world Image Restoration

### v1

@@ 4.1 Experimental Settings
Compared Methods. We compare RDDM against best best-performing traditional one-stage method and two-stage ISPIR models methods, as shown in Table 1. For one-stage methods, we choose JDnDmSR (Xing & Egiazarian, ) and SwinIR (Liang et al., 2021) as our baseline. For two-stage ISPIR methods, we choose PIPNet (A Sharif et al., 2021) as the DN and DM module for ISP and diffusion-based IR methods as our Real-IR baselines.
Implementation Details. We train RDDM with the AdamW optimizer at a learning rate of . The entire training process spans 150000 steps with a batch size of 16. The rank of LoRA in the RVAE Encoder and the diffusion network is set to 4. We employ DAPE as the sRGB domain text prompts extractor.
4.2 Comparisons with State-of-the-Arts
@@
Quantitative Comparisons. The quantitative comparisons among the competing methods on the three test datasets are presented in Table 1. Compared against other approaches, RDDM ranks top 3 in terms of all metrics, including full-reference fidelity metrics like PSNR, SSIM and perceptual quality metrics such as LPIPS and DISTS, on DIV2K-Val, DRealSR and RealSR datasets, except for only PSNR on RealSR. While JDnDmSR and SwinIR achieve slightly higher PSNR and SSIM scores than RDDM on the RealSR dataset, they exhibit markedly inferior performance on other metrics, particularly on no-reference metrics like NIQE, MUSIQ, and CLIPIQA, and distribution alignment metric FID. This suggests that their generative capabilities are substantially weaker than RDDM. In addition, RDDM achieves comparable generative performance with diffusion-based methods, while considerably outperforming them from the aspects of image fidelity metrics such as PSNR and SSIM. Overall, RDDM achieves superior image fidelity and comparable generation capability to the state-of-the-art methods. Note that Table 4.1 further shows the Params and FLOPs of the competing methods. While presenting adequate performace, RDDM has the least Params and FLOPs among diffusion-based models.
@@
Qualitative Comparison. Fig. 5 presents the visual comparisons of RDDM on RealSR, DIV2K-val and DRealSR, along with traditional one-stage and two-stage methods. In our first example, the wall textures generated by JDnDmSR, SwinIR, and DiffBIR are notably blurry, with JDnDmSR exhibiting a significant color deviation. PASD, SeeSR, StableSR, and OSEDiff produce more textures but still fail to capture finer details. SUPIR generates clearer results, yet the textures appear highly unnatural with numerous artifacts. RDDM effectively leverages the detailed information from sensor RAW data to produce more realistic wall textures with higher clarity. The second example draws the same conclusion. It indicates that fully utilizing the abundant yet often lost information in sensor RAW can effectively address the artifact issue prevalent in existing diffusion-based methods. The third and fourth comparisons further manifests that RDDM is capable of generating clear textures with finer details, higher fidelity and less noise. More visualization comparison results can be found in the Appendix. To further investigate the user preferences about these results, we conduct a user study comparing our method on RealSR, DrealSR, and DIV2K-val test datasets, with 5 participants involved. For each set of comparison images, users select their preferred result. As shown in Fig. 6, the results demonstrate that our method significantly outperforms state-of-the-art methods in terms of perceptual quality.
@@
| sRGB VAE | 22.2130 | 0.6200 | 0.3654 | 0.2035 | 58.9168 | 4.4919 | 63.6258 | 0.6912 |
| RVAE | 25.1264 | 0.7092 | 0.2546 | 0.1589 | 36.8671 | 4.1286 | 65.8881 | 0.6723 |
4.3 Ablation Study
@@
| Setting | PSNR | SSIM | LPIPS | FID |
| 1 | 25.2625 | 0.8017 | 0.1719 | 41.8118 |
| 2 | 27.1487 | 0.7035 | 0.3068 | 48.4675 |
| 3 | 27.6892 | 0.7176 | 0.2787 | 38.2007 |
| 4 | 32.5424 | 0.9082 | 0.0533 | 11.6156 |
The importance of RVAE. To illustrate the importance of the proposed RVAE, we substitute it with a pre-trained sRGB VAE. The quantitative comparison on RealSR testset is shown in Table 3. The RDDM employing RVAE outperforms its counterpart utilizing a pre-trained sRGB VAE in both fidelity metrics and perceptual quality metrics with the exception of CLIPIQA. Fig. 7 illustrates that RVAE can significantly mitigate the color deviation issue in RDDM. This is attributed to the fact that the pre-trained sRGB VAE is incapable of adapting to sensor RAW images. In contrast, our RVAE is capable of effectively encoding the sensor RAW data and decoding it into the linear domain images.
@@
Setting of RAW Domain VAE. We systematically evaluate four transfer strategies: (1) We directly use a pre-trained sRGB domain VAE. The model fails to reconstruct the regular Bayer mosaic present in the sensor RAW data and is unable to produce a faithful linear domain image, as depicted in Fig. 8(a). (2) We construct the end-to-end training of the VAE with sensor RAW and linear image pairs. We observe that the model tends to produce images with significant mosaic textures, as illustrated in Fig. 8(b). (3) We first train the encoder and decoder with linear domain images, and then freeze the decoder while training the LoRA layers applied to the encoder with sensor RAW and linear image pairs. However, the result is still dominated by noticeable mosaic textures, as depicted in Fig. 8(c). (4) Following the same first training stage as (3), we freeze the decoder and conduct joint training by incorporating LoRA layers into both the encoder and the diffusion network. As shown in Fig. 8(d), the model successfully recovers fine image details and markedly mitigates color deviation. Table 5 demonstrates that our method attains state-of-the-art reconstruction metrics.
The effectiveness of dual domain loss. We evaluate the contribution of different domain loss in Table 4. Training RDDM with supervision solely in the RAW domain yields satisfactory performance in image generation metrics but performs poorly in both fidelity and perceptual quality metrics, and also introduces localized color deviation issues, as illustrated in Fig. 9(a). In contrast, training RDDM exclusively in the sRGB domain results in sub-optimal performance across all metrics, although it mitigates the color deviation problem, as shown in Fig. 9(b). Training RDDM with supervision in both the RAW and sRGB domains leads to substantial improvements in fidelity and image perceptual quality metrics. However, certain image generation metrics, such as MUSIQ and CLIPIQA, experience a slight decline, while there is a significant improvement in visual quality, as depicted in Fig. 9(c).
The comparison of text prompt extractors. Ultimately, we undertake a systematic evaluation of alternative text prompt extractors. We devise three distinct text extraction strategies. (1) We directly apply a pre-trained sRGB text prompt extractor (sRGBPE) to the sensor RAW images. (2) ISPPE first converts the sensor RAW data to sRGB via an ISP and then extracts prompts with the same pre-trained sRGB extractor. (3) Our proposed RPE employs DDNet for joint denoising and demosaicking, followed by a PTP module that maps the result into the sRGB domain before prompt extraction. As shown in Table 6, RPE reliably extract text prompts, effectively activating the priors encoded in the pre-trained diffusion network and yielding consistent improvements on image fidelity metrics such as PSRN and image generation metrics such as DISTS, FID, NIQE, MUSIQ and CLIPIQA.
@@ 5 Conclusion
We propose RDDM, a novel paradigm for Real-IR, restores directly from the sensor RAW. RDDM exploits the unprocessed, detail-rich signals in sensor RAW data to achieve high fidelity and perceptual quality, thereby alleviating the sub-optimal performance commonly observed in existing Real-IR models that rely on lossy sRGB imagery. Furthermore, RDDM is compatible with diverse Bayer pattern sensor RAW and can be extended to multi-frame input scenarios. The proposed paradigm is applicable to tasks for which the RAW data is accessible, including image and video capture on edge devices. We hope that this work facilitates practical applications and following studies bridging generative modeling, image processing of RAW and image restorations.

### v2

@@ 4.2 Comparisons with State-of-the-Arts
Quantitative Comparisons on synthetic RAW benchmark. Table 1 presents the quantitative comparisons on three synthetic test datasets. RDDM ranks in the top 3 for all metrics, including PSNR, SSIM, LPIPS, DISTS, NIQE, MUSIQ, CLIPIQA, and FID, across DIV2K-Val, DRealSR, and RealSR, except for PSNR on RealSR. JDnDmSR and SwinIR achieve slightly higher PSNR and SSIM on RealSR but significantly underperform on other metrics, particularly NIQE, MUSIQ, CLIPIQA, and FID, indicating weaker generative capabilities. RDDM matches diffusion-based methods in generative performance while outperforming them in image fidelity metrics like PSNR and SSIM. Moreover, RDDM maintains the highest efficiency among diffusion-based competitors in terms of Params and FLOPs as shown in Table 8.
@@
Qualitative Comparison on synthetic RAW benchmark. Visual comparisons on DIV2K-Val and DRealSR demonstrate RDDM’s superiority over existing one-stage and two-stage RAW-domain methods, as shown in Fig.9. While competitors like JDnDmSR and SwinIR suffer from blurriness and color bias, and diffusion-based models (e.g., SeeSR, OSEDiff) often fail to recover fine structures, RDDM reconstructs realistic, high-clarity textures. This confirms that RDDM effectively leverages RAW sensor signals to mitigate typical diffusion artifacts. More visualization comparison results are in the Appendix A.3. To further investigate the user preferences about these results, we conduct a user study on RealSR, DRealSR, and DIV2K-val test datasets, with 5 participants involved. For each set of comparison images, users select their preferred result. As shown in Fig. 8, the results demonstrate that our method significantly outperforms state-of-the-art methods in terms of perceptual quality.
Qualitative Comparison on real RAW benchmark. Visual comparisons on the DND dataset (Fig.10) highlight RDDM’s ability to recover clear and realistic structures on complex surfaces, whereas competing methods often over-smooth textures or fail to suppress noise. As depicted in Fig.11, on the multi-exposure SIDD dataset, RDDM consistently restores clean checkerboard patterns across varying exposure levels, avoiding the color deviations prevalent in one-stage models like MambaIR V2. Furthermore, RDDM exhibits robust generalization to unseen sensor patterns. When evaluated on the RealCapture dataset (RGBG Bayer pattern), RDDM successfully uncovers fine foliage and wall details, as illustrated in Fig.12. In contrast, alternative methods suffer from severe residual noise, grid artifacts, and detail loss, confirming RDDM’s superior cross-sensor adaptability.
4.3 Ablation Study
| Setting | PSNR | SSIM | LPIPS | FID |
| 1 | 25.2625 | 0.8017 | 0.1719 | 41.8118 |
| 2 | 27.1487 | 0.7035 | 0.3068 | 48.4675 |
| 3 | 27.6892 | 0.7176 | 0.2787 | 38.2007 |
| 4 | 26.3147 | 0.6703 | 0.3967 | 53.4004 |
| 5 | 25.4509 | 0.6277 | 0.4533 | 64.2444 |
| 6 | 32.5424 | 0.9082 | 0.0533 | 11.6156 |
The Importance of RVAE. To identify the optimal RVAE architecture and training protocol, we conduct a comprehensive evaluation of six adaptation strategies for RAW-domain reconstruction. The qualitative and quantitative results are summarized as follows:
(i) Baseline Adaptations (Settings 1–2): directly employing a pre-trained sRGB VAE or training from scratch yields either severe degradation or persistent mosaic artifacts (Fig.13a-b).
(ii) LoRA-based Fine-tuning (Settings 3–5): furthermore, fine-tuning only the encoder, decoder, or both via LoRA proves insufficient to bridge the domain gap between sRGB and RAW, as mosaic patterns remain prominent (Fig.13c-e).
(iii) Proposed Joint Training (Setting 6): the optimal strategy involves freezing a pre-trained linear domain decoder while jointly optimizing the encoder and diffusion network via LoRA. This configuration effectively eliminates artifacts, as demonstrated in Fig.13(f).
Quantitative metrics in Table 15 further validate that Setting 6 achieves the highest reconstruction fidelity. Furthermore, End-to-end evaluations within the RDDM framework (Table 2) confirm that our RVAE strategy provides the most robust foundation for high-quality RAW-to-linear restoration, significantly outperforming alternative adaptation protocols. Moreover, Fig.15 confirms that RVAE effectively mitigates chromatic deviations.
| Setting | PSNR | SSIM | LPIPS | DISTS | FID | NIQE | MUSIQ | CLIPIQA |
| 1 | 20.5609 | 0.5797 | 0.3930 | 0.2071 | 73.5659 | 4.0762 | 62.7875 | 0.6431 |
| 2 | 22.3752 | 0.6484 | 0.4785 | 0.3307 | 138.4347 | 6.5993 | 47.2304 | 0.4233 |
| 3 | 21.0398 | 0.6115 | 0.5710 | 0.4235 | 143.3710 | 10.5422 | 37.3569 | 0.3745 |
| 4 | 21.3850 | 0.6080 | 0.4250 | 0.2901 | 108.3618 | 7.5271 | 52.8636 | 0.2931 |
| 5 | 21.3886 | 0.5732 | 0.4911 | 0.3166 | 135.6228 | 7.5537 | 38.7119 | 0.1483 |
| 6 | 25.1264 | 0.7092 | 0.2546 | 0.1589 | 36.8671 | 4.1286 | 65.8881 | 0.6723 |
@@ 4.3 Ablation Study
| All-in-One LoRA | 24.5355 | 0.7003 | 0.2550 | 0.1555 | 40.6767 | 4.1165 | 65.0143 | 0.6660 |
| CMB-LORA | 25.1264 | 0.7092 | 0.2546 | 0.1589 | 36.8671 | 4.1286 | 65.8881 | 0.6723 |
The Necessity of CMB-LoRA. We investigate the parameter-sharing strategy within the CMB-LoRA module by comparing our pattern-specific design against a shared-weight baseline. As reported in Table 3, the shared configuration leads to performance degradation across diverse sensor layouts. This decline stems from the inherent spatial disparities among heterogeneous Bayer patterns (e.g., RGGB, BGGR, and RGBG). While a single weight matrix struggles to capture such diverse local structural characteristics, our pattern-specific design provides the necessary flexibility to model the unique statistical distributions of each RAW format, ensuring robust adaptation across heterogeneous sensors.
@@
The Effectiveness of Dual Domain Loss. We evaluate the impact of domain-specific losses in Table 4. Exclusively RAW-domain training (Setting 1) yields high non-reference scores (e.g., MUSIQ, CLIPIQA) but introduces chromatic artifacts due to the lack of color constraints. Conversely, sRGB-only optimization (Setting 2) ensures color consistency but sacrifices structural fidelity by discarding sensor-level signals. By imposing dual-domain supervision, RDDM achieves an optimal trade-off between structural authenticity and visual realism. As illustrated in Fig.16, while this strategy shows a marginal decline in reference-less metrics, which often over-reward texture intensity, it significantly improves global color stability and alignment with the ground truth, prioritizing holistic fidelity over local high-frequency noise.
@@
The Comparison of Text Prompt Extractors. To evaluate the efficacy of our proposed RPE, we compare it against two alternative prompt-extraction strategies: (1) Direct sRGB-Extractor, which utilizes a pre-trained sRGB model on RAW data; and (2) ISP-based Conversion, which extracts prompts from RAW images pre-processed by a standard ISP. As summarized in Table 5, the proposed RPE consistently outperforms both baselines across all metrics, including reconstruction fidelity (PSNR, DISTS) and perceptual quality (FID, NIQE, MUSIQ, CLIPIQA). By generating more precise prompts, RPE effectively activates the rich generative priors of the diffusion model, facilitating high-fidelity restoration and superior detail synthesis that standard or ISP-processed extractors fail to achieve.
@@ 5 Conclusion
In this paper, we present RDDM, a pioneering paradigm for Real-IR that operates directly from sensor RAW data. By integrating the robust generative priors of the diffusion model, RVAE with a specialized CMB-LoRA module, RDDM effectively addresses the inherent limitations of traditional sRGB-based restoration. Unlike conventional methods that struggle with the information loss incurred by irreversible ISP pipelines, our approach leverages the unprocessed native signal of RAW images to recover superior structural details and textures. While the current 8 down-sampling architecture inherited from Stable Diffusion 2.1 imposes certain constraints on fine-scale fidelity. Future research will focus on developing VAE encode strategies to further preserve sub-pixel spatial information and push the boundaries of high-fidelity Real-IR.
Appendix A Appendix
The appendix provides additional technical details, experimental results, and visual comparisons to support the main paper. Specifically, it includes:
A.1 ISP and Inverse ISP
ISP and Inverse ISP. The RAW sensor data obtained by a camera is fundamentally different from the sRGB images, which closely resemble human visual perception. Indeed, it is necessary to process RAW images through an ISP to obtain the final sRGB images. Fig.17 illustrates the processing pipelines of ISP and InverseISP. We modify the InverseISP of Brooks et al. (2019) and propose a differentiable ISP approach that can map the model’s output to the sRGB domain for optimization. During the ISP process, denoising and demosaicing are ill-posed problems, and we optimize these two tasks jointly with the Real-IR task. Above all, in the ISP process, the Automatic White Balance (AWB) algorithm multiplies the red and blue channels by gains to produce an image that appears to be lit under “neutral” illumination, which can be formulated as Equation 9:
where , , and are the linear-domain image processed by AWB, the linear image processed by denoising and demosaicing, and the pixel-wise channel gain, respectively.
Secondly, the color correction algorithm converts its own “camera space” sRGB color measurements to sRGB values by a color correction matrix as shown in Equation 10:
where , are the linear-domain image processed by the CCM algorithm and the color correction matrix, respectively.
Thirdly, since humans are more sensitive to gradations in the dark areas of images, gamma compression is typically used to allocate more bits of dynamic range to low-intensity pixels as stated in Equation 11:
Note that we set to prevent numerical instability during training.
InverseISP is the inverse process of ISP, where the mosaicing algorithm acquires the RAW image with a CFA by extracting the corresponding pixel values from the three channels. The noise in RAW images mainly comes from two sources: photon arrival statistics (shot noise) and imprecision in the readout circuitry (read noise). We can approximate these two types of noise as a single heteroscedastic Gaussian distribution defined in Equation 12:
The linear-domain image is obtained by Equation 13:
where is the RAW image processed by mosaic algorithm and is the noise intensity added onto to obtain the Sensor RAW . and are the function of ISO light sensitivity level.
A.2 Quantitative Comparisons with GAN-based Methods
| DIV2K-Val | ISP+BSRGAN | 24.4124 | 0.6500 | 0.4174 | 0.2173 | 33.8937 | 5.1222 | 44.2923 | 0.4580 |
| ISP+Real-ESRGAN | 24.2373 | 0.6426 | 0.3888 | 0.2007 | 35.3941 | 4.8658 | 48.6062 | 0.5157 | |
| ISP+LDL | 24.2922 | 0.6243 | 0.4684 | 0.2540 | 34.9855 | 5.5153 | 29.8086 | 0.3714 | |
| ISP+FeMaSR | 23.5889 | 0.6066 | 0.4347 | 0.2273 | 34.9390 | 5.7403 | 43.3786 | 0.5187 | |
| Ours | 23.7416 | 0.6296 | 0.2540 | 0.1197 | 23.8028 | 3.3627 | 65.4202 | 0.6737 | |
| DRealSR | ISP+BSRGAN | 30.1134 | 0.8282 | 0.3196 | 0.1982 | 20.5962 | 5.6748 | 42.4667 | 0.5105 |
| ISP+Real-ESRGAN | 29.1309 | 0.8055 | 0.3361 | 0.2042 | 23.5094 | 5.5440 | 46.0976 | 0.5548 | |
| ISP+LDL | 29.9312 | 0.7967 | 0.3640 | 0.2173 | 20.9155 | 5.9973 | 29.3485 | 0.4047 | |
| ISP+FeMaSR | 28.1176 | 0.7460 | 0.4007 | 0.2356 | 24.3254 | 6.3530 | 41.3755 | 0.5844 | |
| Ours | 28.3495 | 0.7892 | 0.2719 | 0.1649 | 17.4825 | 4.6852 | 57.0696 | 0.7035 | |
| RealSR | ISP+BSRGAN | 27.0999 | 0.7648 | 0.3190 | 0.1984 | 50.7266 | 5.0958 | 50.2340 | 0.5052 |
| ISP+Real-ESRGAN | 26.3844 | 0.7474 | 0.3332 | 0.1995 | 49.3939 | 4.5754 | 55.7670 | 0.5498 | |
| ISP+LDL | 26.5426 | 0.7186 | 0.3871 | 0.2141 | 53.1533 | 5.2573 | 33.9660 | 0.3656 | |
| ISP+FeMaSR | 25.5820 | 0.6898 | 0.3891 | 0.2278 | 53.7582 | 5.8086 | 51.1097 | 0.5917 | |
| Ours | 25.1264 | 0.7092 | 0.2546 | 0.1589 | 36.8671 | 4.1286 | 65.8881 | 0.6723 |
We compare RDDM against four two-stage ISPGAN-based-IR methods. For the first ISP stage, we again use PIPNet A Sharif et al. (2021) as the DN and DM module; for the second IR stage, we use BSRGAN Zhang et al. (2021), Real-ESRGAN Wang et al. (2021), LDL Liang et al. (2022), and FeMaSR Chen et al. (2022) as our baselines. The results are shown in Table 6. Admittedly, two-stage methods involving GAN-based IR models demonstrate better fidelity results as expected, i.e., higher PSNR and SSIM metrics. However, their image generation capability is far behind RDDM. Visualization comparisons are shown in Fig. 18.
A.3 More Qualitative Comparisons
To further demonstrate the perceptual superiority of our approach, we provide extended qualitative comparisons in Fig.19. These comparisons encompass both RAW-domain one-stage methods and representative two-stage pipelines (i.e., ISP followed by sRGB domain IR). The visual results across diverse challenging scenarios consistently showcase that RDDM achieves superior texture restoration and color fidelity, effectively mitigating the artifacts typically found in traditional multi-stage processing.
A.4 Examples of Synthetic RAW Data
Following the data synthesis pipeline detailed in Fig.17, we provide representative examples of our synthesized RAW images in Fig.20. These samples visually demonstrate that our pipeline effectively simulates realistic sensor noise and Bayer mosaics. By bridging the gap between synthetic degradations and real-world RAW characteristics, these generated images provide high-quality supervision, which is fundamental to the robust performance of RDDM.